OpenStudy (anonymous):

Solve for x: −2x − 4 > 8 x < −6 x > −6 x < −2 x > −2

OpenStudy (anonymous):

@gorv can you help? :/

OpenStudy (anonymous):

What happens if you add 4 to both sides?

OpenStudy (anonymous):

Im really not sure how to solve this and I have one other question. I have to finish a few more tests or I'm going to be withdrawn.

OpenStudy (gorv):

-2x-4+4>8+4 -2x>12 -2x-12>0 -12>2x divide both side by 2 -6>x or x<-6

OpenStudy (anonymous):

Thank you! Can you help with one more?

OpenStudy (anonymous):

Solve them like regular equations, just remember if you divide an inequality by a negative, the sign flips.

OpenStudy (anonymous):

Do you know how to solve -2x -4 = 8 ?

OpenStudy (anonymous):

Solve for x: 3 - (2x - 5)<-4(x + 2 ) x < - 8 x > - 8 x < - 3 x > -3

OpenStudy (gorv):

you solve it yourself

OpenStudy (gorv):

open the bracket first and take minus sign inside

OpenStudy (anonymous):

Im really confused on this subject lately, I dont know how to solve this.

OpenStudy (gorv):

try it i m here

OpenStudy (gorv):

-(x-y)=-x-(-y)=-x+y

OpenStudy (anonymous):

That totally just confused me more lol.

OpenStudy (gorv):

ok when two minus sign are their they will become plus and anything outside bracket will multiplied inside after removing bracket

OpenStudy (gorv):

now???

OpenStudy (gorv):

helloooooooooo plzzzzzzzzzzzzzzzzzzzzzzzz type something buddy i m waiting

OpenStudy (anonymous):

Gahh. So 3 * 2x+5?

OpenStudy (anonymous):

Im so confused.

OpenStudy (gorv):

3 - (2x - 5)<-4(x + 2 ) minus sign is there after 3

OpenStudy (gorv):

3-2x+5<-4x-8 8-2x<-4x-8 4x-2x<-8-8 2x<-16 x<-8

OpenStudy (anonymous):

Thank you sososo much!